### v2

@@ 1 Introduction
Drought stress adversely affects the growth, development and yield of plants. The rise in global temperature coupled with the growing demand for water severely depletes soil water reserves. Water deficit and temperature stress can cause damage to the reproductive growth phase of crops, resulting in a substantial reduction in yield [1]. Furthermore, prolonged exposure to water deficit stress particularly in the reproductive stage can significantly affect grain quality. Ensuring sustainable crop yields in the face of escalating drought conditions has become a great challenge [2]. The plant’s response to drought stress is influenced by a number of factors such as stress duration, severity, genotype and its developmental stage [3]. In plants, drought stress in early phase induces reduced stomatal conductance, transpiration and net carbon assimilation, whereas mid- and late-phases result in visible leaf senescence, stagnant growth and wilting. The early detection of drought induced phenotypic changes in plants, facilitates in preventing the adverse impacts of drought from irreversible damage through timely irrigation and other management practices. The point based, contact type and destructive methods based estimation of drought stress impact on crops are limited by good approximation of representative area as well as sample size and hence, difficult to implement over larger agricultural production systems. On the contrary, image-based, non-invasive and high-throughput analyses of subtle phenotypical changes in plants due to drought and other stresses, provides an efficient means of monitoring stress levels in crops and good prediction of drought stress impact [4]. Various imaging techniques and sensors are utilized to precisely capture stress responses, including simple red-blue-green (RGB) , thermal, multispectral, hyperspectral, or fluorescence imaging methods [5]. However, traditional image processing generates significant amount of variation in light intensity, shading, occlusion, etc. leading to variation in quality thereby increases complexity of image processing [6]. Machine learning (ML) techniques, on the other hand, solve such complex problems which are nonlinear in nature by identifying key feature to be extracted from the acquired image in real time, based on knowledge and/or domain expertise of the user. Hence, ML aids in better decision-making and informed actions in real-world scenarios with minimal human intervention [7]. Deep learning (DL) [8] techniques however, help in extracting different information from the raw image data during input training, by means of various convolutions which allows larger learning capabilities, that result in the best classification/regression and higher performance and precision. Convolutional Neural Networks (CNNs) have been developed for providing a much-detailed representation of the image features, for a more discriminative and detailed analysis of plant images, leading to highly accurate classifications [9]. Many studies have used deep CNN models to detect drought stress in affected plants, mostly solved either by finely tuned CNNs or by training from scratch.
@@
Early work by Zhuang et al. [10] utilized a combination of segmentation, color, and texture feature extraction, employing a gradient boosting decision tree (GBDT) method to detect water stress in maize. However, a subsequent study using the same RGB dataset demonstrated that a deep convolutional neural network (DCNN) significantly outperformed GBDT in terms of performance [11]. Sankararao et al. [12] developed a comprehensive pipeline for detecting water stress in groundnut canopies using hyperspectral imaging. The pipeline included steps for image quality assessment, denoising, and band selection, followed by classification using Support Vector Machine (SVM), Random Forest (RF), and Extreme Gradient Boosting (XGBoost/XGB). Transfer learning with a pre-trained DenseNet-121 model has been utilized to classify soybean color images into four levels of drought severity [13]. Similarly, a controlled dataset is created for chickpea crops to identify water stress at different stages—control, young seedling, and pre-flowering—using a CNN-LSTM hybrid model where CNNs that served as feature extractors and LSTM models for predicting water stress levels [14]. The evaluation of various CNN models, including AlexNet, GoogLeNet, and Inception V3, for distinguishing between stress and non-stress conditions in water-sensitive crops such as maize, okra, and soybean, revealed that GoogLeNet achieved the highest identification accuracy [15]. Chlorophyll fluorescence images of wheat canopies have been studied employing a multi-step process involving segmentation and feature extraction with a correlation-based gray-level co-occurrence matrix (CGLCM) and color features, followed by classification using different ML classifiers. The tree-based methods, particularly random forest (RF) and extra trees classifier, is found giving superior performance [16]. A dataset of aerial images of potato canopies has been studied using a deep learning-based model for drought stress identification, leveraging various imaging modalities and their combinations [17]. An explainable deep learning framework was later proposed using the same dataset, based on CNN architectures and transfer learning, achieving improved accuracy [18]. A regression approach is applied to predict the drought tolerance coefficient using hyperspectral images of tea canopies; the comparison of SVM, RF, and Partial Least Squares Regression (PLSR) revealed that SVM was the most effective [19]. Machine learning and deep learning methods were compared, including DNN, SVM, and RF, for drought stress detection using full spectra and first-order derivative spectra [20]. Goyal et al. [21] created a dataset of RGB images of maize crops and proposed a custom-designed CNN model that outperformed existing state-of-the-art CNN architectures in early drought stress detection. However, Vision Transformers (ViTs) offer notable advantages over CNNs in capturing long-range dependencies and global context due to their self-attention mechanism, whereas CNNs rely on local receptive fields and often struggle to capture relationships across various parts of an image [22]. Additionally, ViTs exhibit greater flexibility with input image sizes and handle complex patterns and data variations more effectively [23].
@@
The self-attention mechanism in ViTs enables more accurate classification by considering the entire image context, which enhances accuracy and robustness compared to traditional CNNs [24]. For example, Dosovitskiy et al. demonstrated that ViTs could outperform CNNs in image classification tasks, highlighting their potential for diversified applications [22]. Recent studies have effectively applied ViTs both in customized form [25, 26] and CNN- ViT hybrid form to plant disease identification [27]. Thakur et al. [28] developed a lightweight model that combines convolutional blocks from VGG 16 and Inception V7 with transformer components such as multi-head attention and multi-layer perceptron to effectively identify a wide range of plant diseases across multiple crops. The model leverages the local feature extraction capabilities of CNNs and the global feature modeling strength of vision transformers, enabling simultaneous extraction of both local and global features from images. ViT has outperformed Inception V3 in terms of accuracy when distinguishing among nine different tomato leaf disease classes [29]. Parez et al. fine-tuned the vision transformer with a four-fold reduction in training parameters, achieving higher accuracy compared to CNN models for disease identification across three datasets [30]. Yu et al. [31] proposed a framework that integrates soft split token embedding and depth-wise convolutional modules into the vision transformer architecture, resulting in improved accuracy. Replacing the MLP module in the ViT encoder block with an Inception module improved accuracy in multi-crop disease classification while reducing the number of trainable parameters [32]. Thai et al. optimized the ViT architecture for cassava leaf disease detection by pruning less important attention heads and using sparse matrix operations, achieving a 2% improvement in F1-score along with reduced model size and training costs [33]. Hemalatha et al. developed a plant disease localization and classification model which uses co-scale, co-attention, and cross-attention mechanisms with a vision transformer in a multi-task learning framework [34]. Li et al. integrated a convolutional block attention module into the standard ViT encoder, enabling the network to filter out irrelevant information and focus on essential features, leading to improved crop disease classification in rice, wheat, and coffee [35]. Vallabhajosyula et al. [36] proposed a novel framework that combines a transformer encoder with ResNet9 for plant disease classification, outperforming several classical CNN-based models. From the extensive literature review, we observe that while most existing works focus primarily on improving accuracy, few address the reduction of trainable parameters, and none explore the explainability of transformers. Our proposed work differs in two key aspects: first, by deciphering the attention mechanism, we emphasize model explainability; second, we devised ViT and support vector machine (SVM) combined framework and investigated its performance.
@@ 2.1 Preparing the Data
The potato crop aerial images utilized in this study have been sourced from a publicly accessible dataset that encompasses multiple modalities [37, 17]. Collected from a field at the Aberdeen Research and Extension Center, University of Idaho, these images serve as valuable resources for training machine learning models dedicated to crop health assessment in precision agriculture applications. Acquired using a Parrot Sequoia multi-spectral camera mounted on a 3DR Solo drone, the dataset features an RGB sensor with a resolution of $4,608\times 3,456$ pixels and four monochrome sensors capturing narrow bands of light wavelengths: green (550nm), red (660nm), red-edge (735nm), and near-infrared (790nm), each with a resolution of $1,280\times 960$ pixels. The drone flew over the potato field at a low altitude of 3 meters, with the primary objective of capturing drought stress in Russet Burbank potato plants attributed to premature plant senescence.
@@
The dataset comprises of 360 RGB image patches in JPG format, each measuring 750×750 pixels. These patches were derived from high-resolution aerial images through cropping, rotating, and resizing operations. Data augmentation was applied to an initial set of 300 images as per procedure in Butte et al. [17], expanding the training dataset to 1,500 images. The remaining 60 images were reserved exclusively for testing. No data augmentation was performed on the test images to ensure an unbiased evaluation. Training classification models requires labeled data with annotated regions of interest. In this study, the targets were regions containing healthy and stressed potato plants. These two conditions were visually distinguishable based on color—healthy plants appeared green, whereas stressed plants exhibited a yellowish hue. Manual annotation was performed using the open-source graphical tool LabelImg [38], allowing bounding boxes to be drawn around both healthy and stressed regions. The resulting annotations, including class labels and bounding box coordinates, were saved and used to generate ground truth data for training the proposed models.
@@ 2.2 Vision Transformer (ViT)
A vision transformer (ViT) is a type of neural network architecture that has revolutionized the field of computer vision [39]. Unlike traditional convolutional neural networks (CNNs), which process images pixel by pixel, ViT processes the input images in a sequential manner by dividing them into fixed-size patches, linearly embedding these patches, and then applying self-attention mechanisms for capturing global dependencies [22]. The VIT architecture used for our work is inspired by Dosovitskiy et al.[22] and depicted in Fig 2(a). It processes images by dividing them into fixed-size embedded patches, linearly transforming these patches, and treating them as sequences. The transformer architecture is then applied, incorporating multi-head self-attention mechanisms [40] that enable the model to capture long-range dependencies within the image. Layer normalization is applied before and after the multi-head attention, ensuring stable training, and a Multilayer Perceptron (MLP) head is added to the transformer’s global representation for task-specific processing.
@@ 2.2.1 ViT Architecture
The proposed ViT architecture is based on the layers in [41] and Dosovitskiy et al.[22] and comprises five main components: patch embedding, positional encoding, transformer encoder , normalization layer and a classification head. This is illustrated in Fig 2(a).
@@ 2.3.1 Attention Maps
Attention maps act as a visual representation of the weights assigned by the self-attention mechanism to each patch. These maps can be visualized for each self-attention layer within the ViT model, as each layer progressively learns more intricate relationships between these image features. Higher weights indicate a greater focus on a specific patch and its connection to others. By analyzing these maps, we can essentially see the model’s ”thought process” during image understanding. We can identify which regions it prioritizes for information extraction. The following section highlights the computation behind the attention mechanism [41].
@@
3 Class VisionTransformerBinary($vit\_model$):
       Data: $vit\_model$: Vision Transformer model
@@ 3 Class VisionTransformerBinary($vit\_model$):
5       Initialize $vit$ with $vit\_model$;
6       Initialize $fc$ with a linear layer of appropriate output size;
7       Initialize $attn\_weights$ as an empty list;
@@ 9       Function forward($x$):
10             Clear $attn\_weights$;
@@
12             Initialize $hooks$ as an empty list;
13             for each $block$ in $vit.encoder.layers$ do
@@ 12             Initialize $hooks$ as an empty list;
16                   Append $hook$ to $hooks$;
@@ 20             $x\leftarrow fc(x)$;
21             for each $hook$ in $hooks$ do
22                   Remove $hook$;
23
@@ 28       Function _capture_attn_weights($module,input,output$):
29             Extract $query,key,value$ from $input$;
30             Append $(query,key,value,output)$ to $attn\_weights$;
31
32
@@ 33       Function _get_output_size():
34             Initialize $output$ with a forward pass of zeros through $vit$;
35             return size of the last dimension of $output$;
36
37
38       Function get_attention_weights():
39             return $attn\_weights$;
@@ 2.4.1 Support Vector Machine (SVM)
Support Vector Machines [42] aim to find a hyperplane that best separates a given set of data points into different classes. Given a set of labeled data points $(\mathbf{x}_{1},y_{1}),(\mathbf{x}_{2},y_{2}),\ldots,(\mathbf{x}_{n},y_{n})$ where $\mathbf{x}_{i}$ is the feature vector of the $i$-th data point, and $y_{i}$ is its corresponding class label ($y_{i}\in\{1,0\}$ for binary classification), SVM seeks to find a hyperplane defined by the equation:

### v3

@@ 1 Introduction
Drought stress adversely affects the growth, development and yield of plants. The rise in global temperature coupled with the growing demand for water severely depletes soil water reserves. Water deficit and temperature stress can cause damage to the reproductive growth phase of crops, resulting in a substantial reduction in yield [seleiman_drought_2021]. Furthermore, prolonged exposure to water deficit stress particularly in the reproductive stage can significantly affect grain quality. Ensuring sustainable crop yields in the face of escalating drought conditions has become a great challenge [dinneny_developmental_2019]. The plant’s response to drought stress is influenced by a number of factors such as stress duration, severity, genotype and its developmental stage [yang_response_2021]. In plants, drought stress in early phase induces reduced stomatal conductance, transpiration and net carbon assimilation, whereas mid- and late-phases result in visible leaf senescence, stagnant growth and wilting. The early detection of drought induced phenotypic changes in plants, facilitates in preventing the adverse impacts of drought from irreversible damage through timely irrigation and other management practices. The point based, contact type and destructive methods based estimation of drought stress impact on crops are limited by good approximation of representative area as well as sample size and hence, difficult to implement over larger agricultural production systems. On the contrary, image-based, non-invasive and high-throughput analyses of subtle phenotypical changes in plants due to drought and other stresses, provides an efficient means of monitoring stress levels in crops and good prediction of drought stress impact [furbank_phenomics_2011]. Various imaging techniques and sensors are utilized to precisely capture stress responses, including simple red-blue-green (RGB) , thermal, multispectral, hyperspectral, or fluorescence imaging methods [Tamimi_2022]. However, traditional image processing generates significant amount of variation in light intensity, shading, occlusion, etc. leading to variation in quality thereby increases complexity of image processing [singh_deep_2018]. Machine learning (ML) techniques, on the other hand, solve such complex problems which are nonlinear in nature by identifying key feature to be extracted from the acquired image in real time, based on knowledge and/or domain expertise of the user. Hence, ML aids in better decision-making and informed actions in real-world scenarios with minimal human intervention [liakos_machine_2018]. Deep learning (DL) [arya_deep_2022] techniques however, help in extracting different information from the raw image data during input training, by means of various convolutions which allows larger learning capabilities, that result in the best classification/regression and higher performance and precision. Convolutional Neural Networks (CNNs) have been developed for providing a much-detailed representation of the image features, for a more discriminative and detailed analysis of plant images, leading to highly accurate classifications [jiang_convolutional_2020]. Many studies have used deep CNN models to detect drought stress in affected plants, mostly solved either by finely tuned CNNs or by training from scratch.
@@
Early work by Zhuang et al. [zhuang_early_2017] utilized a combination of segmentation, color, and texture feature extraction, employing a gradient boosting decision tree (GBDT) method to detect water stress in maize. However, a subsequent study using the same RGB dataset demonstrated that a deep convolutional neural network (DCNN) significantly outperformed GBDT in terms of performance [an_identification_2019]. Sankararao et al. [sankararao_machine_2023] developed a comprehensive pipeline for detecting water stress in groundnut canopies using hyperspectral imaging. The pipeline included steps for image quality assessment, denoising, and band selection, followed by classification using Support Vector Machine (SVM), Random Forest (RF), and Extreme Gradient Boosting (XGBoost/XGB). Transfer learning with a pre-trained DenseNet-121 model has been utilized to classify soybean color images into four levels of drought severity [ramos-giraldo_drought_2020]. Similarly, a controlled dataset is created for chickpea crops to identify water stress at different stages—control, young seedling, and pre-flowering—using a CNN-LSTM hybrid model where CNNs that served as feature extractors and LSTM models for predicting water stress levels [azimi_intelligent_2021]. The evaluation of various CNN models, including AlexNet, GoogLeNet, and Inception V3, for distinguishing between stress and non-stress conditions in water-sensitive crops such as maize, okra, and soybean, revealed that GoogLeNet achieved the highest identification accuracy [chandel_identifying_2021]. Chlorophyll fluorescence images of wheat canopies have been studied employing a multi-step process involving segmentation and feature extraction with a correlation-based gray-level co-occurrence matrix (CGLCM) and color features, followed by classification using different ML classifiers. The tree-based methods, particularly random forest (RF) and extra trees classifier, is found giving superior performance [gupta_drought_2023]. A dataset of aerial images of potato canopies has been studied using a deep learning-based model for drought stress identification, leveraging various imaging modalities and their combinations [butte_potato_2021]. An explainable deep learning framework was later proposed using the same dataset, based on CNN architectures and transfer learning, achieving improved accuracy [patra_explainable_2024]. A regression approach is applied to predict the drought tolerance coefficient using hyperspectral images of tea canopies; the comparison of SVM, RF, and Partial Least Squares Regression (PLSR) revealed that SVM was the most effective [chen_hyperspectral_2022]. Machine learning and deep learning methods were compared, including DNN, SVM, and RF, for drought stress detection using full spectra and first-order derivative spectra [dao_plant_2021]. Goyal et al. [goyal_deep_2024] created a dataset of RGB images of maize crops and proposed a custom-designed CNN model that outperformed existing state-of-the-art CNN architectures in early drought stress detection. However, Vision Transformers (ViTs) offer notable advantages over CNNs in capturing long-range dependencies and global context due to their self-attention mechanism, whereas CNNs rely on local receptive fields and often struggle to capture relationships across various parts of an image [dosovitskiy_image_2021]. Additionally, ViTs exhibit greater flexibility with input image sizes and handle complex patterns and data variations more effectively [chen_transunet_2021].
@@
The self-attention mechanism in ViTs enables more accurate classification by considering the entire image context, which enhances accuracy and robustness compared to traditional CNNs [liu_swin_2021]. For example, Dosovitskiy et al. demonstrated that ViTs could outperform CNNs in image classification tasks, highlighting their potential for diversified applications [dosovitskiy_image_2021]. Recent studies have effectively applied ViTs both in customized form [bhowmik_customised_2024, singh_effective_2024] and CNN- ViT hybrid form to plant disease identification [borhani_deep_2022]. Thakur et al. [thakur_vision_2023] developed a lightweight model that combines convolutional blocks from VGG 16 and Inception V7 with transformer components such as multi-head attention and multi-layer perceptron to effectively identify a wide range of plant diseases across multiple crops. The model leverages the local feature extraction capabilities of CNNs and the global feature modeling strength of vision transformers, enabling simultaneous extraction of both local and global features from images. ViT has outperformed Inception V3 in terms of accuracy when distinguishing among nine different tomato leaf disease classes [barman_vit-smartagri_2024]. Parez et al. fine-tuned the vision transformer with a four-fold reduction in training parameters, achieving higher accuracy compared to CNN models for disease identification across three datasets [parez_visual_2023]. Yu et al. [yu_inception_2023] proposed a framework that integrates soft split token embedding and depth-wise convolutional modules into the vision transformer architecture, resulting in improved accuracy. Replacing the MLP module in the ViT encoder block with an Inception module improved accuracy in multi-crop disease classification while reducing the number of trainable parameters [gole_trincnet_2023]. Thai et al. optimized the ViT architecture for cassava leaf disease detection by pruning less important attention heads and using sparse matrix operations, achieving a 2% improvement in F1-score along with reduced model size and training costs [thai_formerleaf_2023]. Hemalatha et al. developed a plant disease localization and classification model which uses co-scale, co-attention, and cross-attention mechanisms with a vision transformer in a multi-task learning framework [hemalatha_multitask_2024]. Li et al. integrated a convolutional block attention module into the standard ViT encoder, enabling the network to filter out irrelevant information and focus on essential features, leading to improved crop disease classification in rice, wheat, and coffee [li_pmvt_2023]. Vallabhajosyula et al. [vallabhajosyula_novel_2024] proposed a novel framework that combines a transformer encoder with ResNet9 for plant disease classification, outperforming several classical CNN-based models. From the extensive literature review, we observe that while most existing works focus primarily on improving accuracy, few address the reduction of trainable parameters, and none explore the explainability of transformers. Our proposed work differs in two key aspects: first, by deciphering the attention mechanism, we emphasize model explainability; second, we devised ViT and support vector machine (SVM) combined framework and investigated its performance.
@@ 2.1 Preparing the Data
The potato crop aerial images utilized in this study have been sourced from a publicly accessible dataset that encompasses multiple modalities [potato_data, butte_potato_2021]. Collected from a field at the Aberdeen Research and Extension Center, University of Idaho, these images serve as valuable resources for training machine learning models dedicated to crop health assessment in precision agriculture applications. Acquired using a Parrot Sequoia multi-spectral camera mounted on a 3DR Solo drone, the dataset features an RGB sensor with a resolution of $4,608\times 3,456$ pixels and four monochrome sensors capturing narrow bands of light wavelengths: green (550nm), red (660nm), red-edge (735nm), and near-infrared (790nm), each with a resolution of $1,280\times 960$ pixels. The drone flew over the potato field at a low altitude of 3 meters, with the primary objective of capturing drought stress in Russet Burbank potato plants attributed to premature plant senescence.
@@
The dataset comprises of 360 RGB image patches in JPG format, each measuring 750×750 pixels. These patches were derived from high-resolution aerial images through cropping, rotating, and resizing operations. Data augmentation was applied to an initial set of 300 images as per procedure in Butte et al. [butte_potato_2021], expanding the training dataset to 1,500 images. The remaining 60 images were reserved exclusively for testing. No data augmentation was performed on the test images to ensure an unbiased evaluation. Training classification models requires labeled data with annotated regions of interest. In this study, the targets were regions containing healthy and stressed potato plants. These two conditions were visually distinguishable based on color—healthy plants appeared green, whereas stressed plants exhibited a yellowish hue. Manual annotation was performed using the open-source graphical tool LabelImg [noauthor_labelimg_nodate], allowing bounding boxes to be drawn around both healthy and stressed regions. The resulting annotations, including class labels and bounding box coordinates, were saved and used to generate ground truth data for training the proposed models.
@@ 2.2 Vision Transformer (ViT)
A vision transformer (ViT) is a type of neural network architecture that has revolutionized the field of computer vision [han_survey_2023]. Unlike traditional convolutional neural networks (CNNs), which process images pixel by pixel, ViT processes the input images in a sequential manner by dividing them into fixed-size patches, linearly embedding these patches, and then applying self-attention mechanisms for capturing global dependencies [dosovitskiy_image_2021]. The VIT architecture used for our work is inspired by Dosovitskiy et al.[dosovitskiy_image_2021] and depicted in Fig 2(a). It processes images by dividing them into fixed-size embedded patches, linearly transforming these patches, and treating them as sequences. The transformer architecture is then applied, incorporating multi-head self-attention mechanisms [parikh_decomposable_2016] that enable the model to capture long-range dependencies within the image. Layer normalization is applied before and after the multi-head attention, ensuring stable training, and a Multilayer Perceptron (MLP) head is added to the transformer’s global representation for task-specific processing.
@@ 2.2.1 ViT Architecture
The proposed ViT architecture is based on the layers in [vaswani_attention_2023] and Dosovitskiy et al.[dosovitskiy_image_2021] and comprises five main components: patch embedding, positional encoding, transformer encoder , normalization layer and a classification head. This is illustrated in Fig 2(a).
@@ 2.3.1 Attention Maps
Attention maps act as a visual representation of the weights assigned by the self-attention mechanism to each patch. These maps can be visualized for each self-attention layer within the ViT model, as each layer progressively learns more intricate relationships between these image features. Higher weights indicate a greater focus on a specific patch and its connection to others. By analyzing these maps, we can essentially see the model’s ”thought process” during image understanding. We can identify which regions it prioritizes for information extraction. The following section highlights the computation behind the attention mechanism [vaswani_attention_2023].
@@
3 Class VisionTransformerBinary(vit_modelvit\_model):
 Data: vit_modelvit\_model: Vision Transformer model
@@ 3 Class VisionTransformerBinary(vit_modelvit\_model):
5   Initialize vitvit with vit_modelvit\_model;
6   Initialize fcfc with a linear layer of appropriate output size;
7   Initialize attn_weightsattn\_weights as an empty list;
@@ 9 Function forward($x$):
10      Clear attn_weightsattn\_weights;
@@
12      Initialize hookshooks as an empty list;
13    for each blockblock in $vit.encoder.layers$ do
@@ 12      Initialize hookshooks as an empty list;
16         Append hookhook to hookshooks;
@@ 21    $x\leftarrow fc(x)$;
22    for each hookhook in hookshooks do
23         Remove hookhook;
@@ 31 Function _capture_attn_weights($module,input,output$):
32      Extract $query,key,value$ from inputinput;
33      Append $(query,key,value,output)$ to attn_weightsattn\_weights;
34
35
36
@@ 37 Function _get_output_size():
38      Initialize outputoutput with a forward pass of zeros through vitvit;
39    return size of the last dimension of outputoutput;
@@ 38      Initialize outputoutput with a forward pass of zeros through vitvit;
42
43 Function get_attention_weights():
44    return attn_weightsattn\_weights;
45
46
47
48
@@ 2.4.1 Support Vector Machine (SVM)
Support Vector Machines [hearst_support_1998] aim to find a hyperplane that best separates a given set of data points into different classes. Given a set of labeled data points $(\mathbf{x}_{1},y_{1}),(\mathbf{x}_{2},y_{2}),\ldots,(\mathbf{x}_{n},y_{n})$ where $\mathbf{x}_{i}$ is the feature vector of the $i$-th data point, and $y_{i}$ is its corresponding class label ($y_{i}\in\{1,0\}$ for binary classification), SVM seeks to find a hyperplane defined by the equation: